OpenStudy (anonymous):

f'(x) = 2x - e^x = 0

OpenStudy (anonymous):

what do you want to find?

OpenStudy (anonymous):

to solve for x?

OpenStudy (anonymous):

i want to find x

OpenStudy (anonymous):

I believe this is just part of the whole question, what actually are you working on?

OpenStudy (anonymous):

Finding the absolute maximum and minimum values of f(x) = 2x - e^x on [0,1]. i'm stuck

OpenStudy (anonymous):

ok, good, then the information given above is not correct f(x) = 2x - e^x f'(x) = 2 - e^x

OpenStudy (anonymous):

You can solve your f'(x) = 0 easily now

OpenStudy (anonymous):

i see that e^x should be 2 but wat about x

OpenStudy (anonymous):

e^x = 2 Take log on both sides: x = log(2), log = natural logarithm

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I believe you know how to determine the absolute extrema

OpenStudy (anonymous):

plugging that into the fucntion with the endpoint values and determining the max and min

OpenStudy (anonymous):

very good, yup

OpenStudy (anonymous):

thx for ur help

OpenStudy (anonymous):

you're welcome